Find the surface which is orthogonal to the one parameter system z = c xy( x2 + y2 )

which passes through the hyperbola , x2 - y2 = a2 , z =0.

1
Expert's answer
2021-06-03T15:37:48-0400

"𝑓(𝑥, 𝑦, 𝑧) =\dfrac{z}{xy(x^2+y^2)}=c"

"f_x=-\dfrac{z(3x^2+y^2)}{x^2y(x^2+y^2)^2}"


"f_y=-\dfrac{z(3y^2+x^2)}{xy^2(x^2+y^2)^2}"

"f_z=\dfrac{1}{xy(x^2+y^2)}"

Let "z(x,y)" is a surface orthogonal to the given system. Then:


"(f_x, f_y, f_z)\cdot(z_x, z_y, -1)=0"

So, we have differential equation:


"-\dfrac{z(3x^2+y^2)}{x^2y(x^2+y^2)^2}z_x-\dfrac{z(3y^2+x^2)}{xy^2(x^2+y^2)^2}z_y"

"-\dfrac{1}{xy(x^2+y^2)}=0"

Divide by "-\dfrac{z}{xy(x^2+y^2)}"


"\dfrac{3x^2+y^2}{x(x^2+y^2)}z_x+\dfrac{3y^2+x^2}{y(x^2+y^2)}z_y+\dfrac{1}{z}=0"

The auxiliary equations:


"\dfrac{x(x^2+y^2)}{3x^2+y^2}dx=\dfrac{y(x^2+y^2)}{3y^2+x^2}dy=-zdz"

Adding first and second and equating to third:


"\dfrac{(x^2+y^2)(xdx+ydy)}{4x^2+4y^2}=-zdz"

"\dfrac{(xdx+ydy)}{4}=-zdz"

giving 


"x^2+y^2+4z^2=c_1"

Subtracting the second equation from the first one:


"\dfrac{(x^2+y^2)(xdx-ydy)}{2(x^2-y^2)}"

"\dfrac{(c_1-4z^2)(xdx-ydy)}{2(x^2-y^2)}=-zdz"


"x^2-y^2=c_2\sqrt{c_1-4z^2}"

"c_2=\dfrac{x^2-y^2}{\sqrt{c_1-4z^2}}"

Using the given conditions "x^2-y^2=a, z=0:"


"a^2=c_1c_2^2"

Then


"a^2=(x^2+y^2+4z^2)\dfrac{(x^2-y^2)^2}{x^2+y^2}"

"x^2+y^2+4z^2=\dfrac{a^2(x^2+y^2)}{(x^2-y^2)^2}"

"z=\sqrt{\dfrac{a^2(x^2+y^2)}{4(x^2-y^2)^2}-\dfrac{x^2+y^2}{4}}"